# Joint multi-dimensional dynamic attention and transformer for general image restoration

### v1

@@ II-A Image Restoration
Image restoration is a long-standing low-level task, in which clean images are restored from degraded ones with noise, rain streak, or blur, etc. The image restoration methods can be categorized as two classes. One is customized image restoration methods, including image deraining [18, 19, 20], dehazing [21, 22, 23], denoising [24, 25], deblurring [26, 27], which are designed for specific degradation. For example, rain streak binary map and rain accumulation map are utilized to indicate rain streak location and appearance prior, respectively [18]. The hazy image is decomposed into three latent layers through three branches and then reconstructed into a clean image in an unsupervised manner[23].
The other is general image restoration methods for various image degradation. MPRNet [28], SwinIR [29], MAXIM [30] are three representative newly proposed methods of CNN-based, transformer-based, and multi-layer perceptron (MLP)-based image restoration methods. Recently, to reduce the complexity of transformer, Uformer [11] and KiT [31] with different local-window self-attention strategies were proposed. Restormer [10] was designed with transposed attention and gated-dconv feed-forward network. In addition, all-in-one approaches [32, 33] that could address degraded images with several common types by one model have emerged.
@@ III-B Multi-dimensional Dynamic Attention Block
| Type | Regular | CondConv/ DyConv | ODConv | MDConv |
| Spatial | ✗ | ✗ | ✓ | ✓ |
| Channel | ✗ | ✗ | ✓ | ✓ |
| Filter | ✗ | ✗ | ✓ | ✓ |
| Kernel | ✗ | ✓ | ✓ | ✗ |
[…]
@@ IV-A Experiment Settings
| Method | Venue | Test100[48] | Rain100H[18] | Rain100L[18] | Test2800[47] | Test1200[50] | Average | FLOPs(G)/ #Params(M) | ||||||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |||
| DerainNet [67] | 17’TIP | 22.77 | 0.810 | 14.92 | 0.592 | 27.03 | 0.884 | 24.31 | 0.861 | 23.38 | 0.835 | 22.48 | 0.796 | 88.29/0.75 |
| RESCAN [68] | 18’ECCV | 25.00 | 0.835 | 26.36 | 0.786 | 29.80 | 0.881 | 31.29 | 0.904 | 30.51 | 0.882 | 28.59 | 0.857 | 32.32/0.15 |
| UMRL [69] | 19’CVPR | 24.41 | 0.829 | 26.01 | 0.832 | 29.18 | 0.923 | 29.97 | 0.905 | 30.55 | 0.910 | 28.02 | 0.880 | 16.50/0.98 |
| PreNet [70] | 19’CVPR | 24.81 | 0.851 | 26.77 | 0.858 | 32.44 | 0.950 | 31.75 | 0.916 | 31.36 | 0.911 | 29.42 | 0.897 | 66.58/0.17 |
| MSPFN [71] | 20’CVPR | 27.50 | 0.876 | 28.66 | 0.860 | 32.40 | 0.933 | 32.82 | 0.930 | 32.39 | 0.916 | 30.75 | 0.903 | 708.44/21.00 |
[…]
| HINet [65] | 21’CVPRW | 30.26 | 0.905 | 30.63 | 0.893 | 37.20 | 0.969 | 33.87 | 0.940 | 33.01 | 0.918 | 33.00 | 0.925 | 170.73/88.67 |
| KiT [31] | 22’CVPR | 30.26 | 0.904 | 30.47 | 0.896 | 36.65 | 0.970 | 33.85 | 0.941 | 32.81 | 0.918 | 32.81 | 0.926 | 43.08/20.60 |
| MAXIM-2S [30] | 22’CVPR | 31.17 | 0.922 | 30.81 | 0.903 | 38.06 | 0.976 | 33.80 | 0.943 | 32.37 | 0.921 | 33.24 | 0.933 | 216.00/14.10 |
| Restormer [10] | 22’CVPR | 32.00 | 0.923 | 31.46 | 0.904 | 38.99 | 0.978 | 34.18 | 0.944 | 33.19 | 0.926 | 33.96 | 0.935 | 140.99/26.13 |
| SFNet [72] | 23’ICLR | 31.47 | 0.919 | 31.90 | 0.903 | 38.21 | 0.974 | 33.69 | 0.937 | 32.55 | 0.911 | 33.56 | 0.929 | 125.43/13.27 |
| MambaIR [73] | 24’ECCV | 32.11 | 0.924 | 32.03 | 0.905 | 38.46 | 0.976 | 34.15 | 0.945 | 33.53 | 0.930 | 34.06 | 0.936 | 229.80/26.78 |
@@
| Method | Venue | Raindrop-A[51] | Raindrop-B[51] | FLOPs(G)/ #Params(M) | ||
| PSNR | SSIM | PSNR | SSIM | |||
| pix2pix [74] | 17’CVPR | 28.02 | 0.855 | 23.50 | 0.715 | 18.16/54.40 |
| AttentiveGAN [51] | 18’CVPR | 31.55 | 0.902 | 24.92 | 0.809 | 89.67/6.24 |
| DuRN [4] | 19’CVPR | 31.24 | 0.926 | 25.32 | 0.817 | 89.39/12.86 |
| IDT[75] | 22’TPAMI | 31.87 | 0.931 | - | - | 61.90/16.00 |
| MAXIM-2S[30] | 22’CVPR | 31.87 | 0.935 | 25.74 | 0.827 | 216.00/14.10 |
| UDR-S2Former[76] | 23’CVPR | 32.64 | 0.943 | - | - | 21.58/8.53 |
@@ IV-B Image Deraining Results
| Method | Venue | GoPro[52] | HIDE[53] | RealBlur-R[54] | RealBlur-J[54] | FLOPs(G)/ #Params(M) | ||||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |||
| Nah et al. [52] | 17’CVPR | 29.08 | 0.914 | 25.73 | 0.874 | 32.51 | 0.841 | 27.87 | 0.827 | 336.00/11.70 |
| DeblurGAN [26] | 18’CVPR | 28.70 | 0.858 | 24.51 | 0.871 | 33.79 | 0.903 | 27.97 | 0.834 | 16.96/14.00 |
| DeblurGAN-v2 [77] | 19’ICCV | 29.55 | 0.934 | 26.61 | 0.875 | 35.26 | 0.944 | 28.70 | 0.866 | 42.00/60.93 |
| DMPHN [78] | 19’CVPR | 31.20 | 0.940 | 29.09 | 0.924 | 35.70 | 0.948 | 28.42 | 0.860 | 235.00/21.70 |
| DBGAN [79] | 20’CVPR | 31.10 | 0.942 | 28.94 | 0.915 | 33.78 | 0.909 | 24.93 | 0.745 | 759.85/11.59 |
| MTRNN [80] | 20’ECCV | 31.15 | 0.945 | 29.15 | 0.918 | 35.79 | 0.951 | 28.44 | 0.862 | 13.72/2.64 |
| MIMO-UNet+ [27] | 21’ICCV | 32.45 | 0.957 | 29.99 | 0.930 | 35.54 | 0.947 | 27.63 | 0.837 | 154.41/16.11 |
| Stripformer [81] | 22’ECCV | 33.08 | 0.962 | 31.03 | 0.940 | - | - | - | - | 177.43/19.71 |
| DDANet [82] | 23’AAAI | 33.07 | 0.962 | 30.64 | 0.937 | 35.81 | 0.951 | - | - | 153.51/16.18 |
[…]
| NAFNet [46] | 22’ECCV | 33.08 | 0.963 | 31.22 | 0.943 | 36.14 | 0.955 | 28.43 | 0.860 | 63.33/67.89 |
| Restormer [10] | 22’CVPR | 32.92 | 0.961 | 31.22 | 0.942 | 36.19 | 0.957 | 28.96 | 0.879 | 140.99/26.11 |
| MambaIR [73] | 24’ECCV | 33.27 | 0.963 | 31.11 | 0.941 | 36.15 | 0.955 | 28.99 | 0.878 | 229.80/26.78 |
@@ IV-C Image Deblurring Results
| Dataset | Evaluation | DnCNN | RIDNet | ADFNet | MIRNet | MPRNet | Uformer-B | MAXIM-3S | Restormer | NAFNet | MambaIR | MDDA-former |
| [2] | [84] | [85] | [86] | [28] | [11] | [30] | [10] | [46] | [73] | |||
| SIDD [55] | PSNR SSIM | 30.78 0.801 | 38.71 0.951 | 39.63 0.958 | 39.72 0.959 | 39.71 0.958 | 39.89 0.960 | 39.96 0.960 | 40.02 0.960 | 40.30 0.962 | 39.89 0.960 | 39.96 0.960 |
| DND [56] | PSNR SSIM | 32.43 0.790 | 39.26 0.953 | 39.89 0.960 | 39.88 0.956 | 39.80 0.954 | 39.98 0.955 | 39.84 0.954 | 40.03 0.956 | - - | 40.04 0.956 | 40.05 0.956 |
| Venue | 17’TIP | 19’ICCV | 23’AAAI | 20’ECCV | 21’CVPR | 22’CVPR | 22’CVPR | 22’CVPR | 22’ECCV | 24’ECCV | - | |
| FLOPs (G) #Params (M) | 43.94 0.67 | 98.13 1.50 | 55.64 7.65 | 786.43 31.79 | 148.55 3.64 | 89.46 50.88 | 320.32 20.60 | 140.99 26.11 | 63.33 67.89 | 229.80 26.78 | 67.38 25.92 |
@@ IV-D Image Denoising Results
| Method | Venue | SOTS-Indoor[57] | SOTS-Outdoor[57] | SOTS-Mix[58] | Haze4K[59] | FLOPs (G)/ #Params (M) | ||||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |||
| DCP [87] | 09’CVPR | 16.62 | 0.818 | 19.14 | 0.861 | 17.88 | 0.816 | 14.01 | 0.760 | - |
| DehazeNet [21] | 16’TIP | 19.82 | 0.821 | 24.75 | 0.927 | 21.02 | 0.870 | 19.12 | 0.840 | 0.58/0.009 |
| MSCNN [88] | 16’ECCV | 19.84 | 0.833 | 22.06 | 0.908 | 20.31 | 0.863 | 14.01 | 0.510 | 0.53/0.008 |
| AOD-Net [89] | 17’ICCV | 20.51 | 0.816 | 24.14 | 0.920 | 20.27 | 0.855 | 17.15 | 0.830 | 0.12/0.002 |
| GridDehazeNet [90] | 19’ICCV | 32.16 | 0.984 | 30.86 | 0.982 | 25.86 | 0.944 | 23.29 | 0.930 | 21.49/0.96 |
| MSBDN [91] | 20’CVPR | 33.67 | 0.985 | 33.48 | 0.982 | 28.56 | 0.966 | 22.99 | 0.850 | 41.54/31.35 |
| FFA-Net [6] | 20’AAAI | 36.39 | 0.989 | 33.57 | 0.984 | 29.96 | 0.973 | 26.96 | 0.950 | 287.80/4.46 |
| AECR-Net [22] | 21’CVPR | 37.17 | 0.990 | - | - | 28.52 | 0.964 | - | - | 52.20/2.61 |
| DeHamer [92] | 22’CVPR | 36.63 | 0.988 | 35.18 | 0.986 | - | - | - | - | 48.93/132.5 |
| DehazeFormer-M [93] | 23’TIP | 38.46 | 0.994 | 34.29 | 0.983 | 30.89 | 0.977 | - | - | 48.64/4.63 |
| DehazeFormer-L [93] | 23’TIP | 40.05 | 0.996 | - | - | - | - | - | - | 279.70/25.44 |
| MITNet [94] | 23’ACM MM | 40.23 | 0.992 | 35.18 | 0.988 | - | - | - | - | 16.42/2.73 |
[…]
| MambaIR [73] | 24’ECCV | 40.58 | 0.995 | 35.41 | 0.986 | 31.24 | 0.979 | 33.17 | 0.987 | 229.80/26.78 |
@@
| Method | Venue | LOL-v1 [60] | LOL-v2-real [61] | FLOPs (G)/ #Params (M) | ||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||
| Retinex-Net [60] | 18’BMVC | 16.77 | 0.560 | 0.381 | 18.37 | 0.723 | 0.425 | 587.47/0.84 |
| KinD [95] | 19’ACM MM | 20.86 | 0.790 | 0.170 | 19.74 | 0.761 | 0.362 | 34.99/8.16 |
| Zero-DCE [96] | 20’CVPR | 14.83 | 0.531 | 0.335 | 14.32 | 0.511 | 0.483 | 5.17/0.08 |
| KinD++ [97] | 21’IJCV | 21.30 | 0.821 | 0.160 | - | - | - | 743.95/8.28 |
| EnlightenGAN [98] | 21’TIP | 17.48 | 0.650 | 0.322 | 18.23 | 0.617 | 0.421 | 61.01/8.64 |
| Sparse [61] | 21’TIP | 17.20 | 0.640 | - | 20.06 | 0.816 | - | 53.26/2.33 |
| DRBN [99] | 21’TIP | 19.55 | 0.746 | 0.316 | 20.13 | 0.820 | 0.304 | 48.61/5.27 |
| URetinex-Net [100] | 22’CVPR | 21.33 | 0.835 | 0.122 | - | - | - | 56.93/0.40 |
| UHDFour [101] | 23’ICLR | 23.09 | 0.870 | - | 21.78 | 0.870 | - | -/17.54 |
[…]
| Restormer [10] | 22’CVPR | 22.43 | 0.823 | 0.141 | 19.94 | 0.827 | 0.292 | 140.99/26.13 |
| MAXIM-2S [30] | 22’CVPR | 23.43 | 0.863 | - | - | - | - | 216.00/14.10 |
| MambaIR [73] | 24’ECCV | 24.05 | 0.857 | 0.113 | 22.11 | 0.834 | 0.262 | 229.80/26.78 |

### v2

@@ II-A Image Restoration
Image restoration is a long-standing low-level task, in which clean images are restored from degraded ones with noise, rain streak, or blur, etc. The image restoration methods can be categorized as two classes. One is customized image restoration methods, including image deraining [18, 19], dehazing [20, 21, 22], denoising [23, 24, 25], deblurring [26, 27], and super-resolution [28, 29], which are designed for specific degradation. For example, rain streak binary map and rain accumulation map are utilized to indicate rain streak location and appearance prior, respectively [18]. Another line of research focuses on general image restoration[30, 31, 32, 10, 11, 33], aiming to handle multiple types of image degradation within a single framework. Representative methods include MPRNet [30], SwinIR [31], and MAXIM [32], which exemplify recent advances based on CNNs, Transformers, and MLPs, respectively. To address the high computational cost of standard Transformers, several efficient architectures have been proposed, such as Uformer [11] and KiT [34], which adopt localized window-based self-attention mechanisms. Restormer [10] further improves efficiency through transposed attention and gated-dconv feed-forward network.
More recently, unified image restoration models [35, 36, 37, 38, 39] capable of addressing multiple degradations (e.g., noise, blur, rain, haze) with a single network have become increasingly popular. For example, GridFormer [37] improves performance in adverse weather by combining residual dense connections with a grid-structured Transformer, facilitating hierarchical feature learning and long-range dependency modeling.
@@ III-B Multi-dimensional Dynamic Attention Block
Type | Regular |
ODConv | MDConv | ||
Spatial | ✗ | ✗ | ✓ | ✓ | ||
Channel | ✗ | ✗ | ✓ | ✓ | ||
Filter | ✗ | ✗ | ✓ | ✓ | ||
Kernel | ✗ | ✓ | ✓ | ✗ | ||
Light-weight | ✓ | ✓ | ✗ | ✓ |
@@ IV-A Experiment Settings
Method | Venue | Test100[54] | Rain100H[18] | Rain100L[18] | Test2800[53] | Test1200[56] | Average |
PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |||||
DerainNet [73] | 17’TIP | 22.77 | 0.810 | 14.92 | 0.592 | 27.03 | 0.884 | 24.31 | 0.861 | 23.38 | 0.835 | 22.48 | 0.796 | 88.29/0.75 | ||
RESCAN [74] | 18’ECCV | 25.00 | 0.835 | 26.36 | 0.786 | 29.80 | 0.881 | 31.29 | 0.904 | 30.51 | 0.882 | 28.59 | 0.857 | 32.32/0.15 | ||
UMRL [75] | 19’CVPR | 24.41 | 0.829 | 26.01 | 0.832 | 29.18 | 0.923 | 29.97 | 0.905 | 30.55 | 0.910 | 28.02 | 0.880 | 16.50/0.98 | ||
PreNet [76] | 19’CVPR | 24.81 | 0.851 | 26.77 | 0.858 | 32.44 | 0.950 | 31.75 | 0.916 | 31.36 | 0.911 | 29.42 | 0.897 | 66.58/0.17 | ||
MSPFN [77] | 20’CVPR | 27.50 | 0.876 | 28.66 | 0.860 | 32.40 | 0.933 | 32.82 | 0.930 | 32.39 | 0.916 | 30.75 | 0.903 | 708.44/21.00 | ||
MPRNet [30] | 21’CVPR | 30.27 | 0.897 | 30.41 | 0.890 | 36.40 | 0.965 | 33.64 | 0.938 | 32.91 | 0.916 | 32.73 | 0.921 | 148.55/3.64 | ||
HINet [71] | 21’CVPRW | 30.26 | 0.905 | 30.63 | 0.893 | 37.20 | 0.969 | 33.87 | 0.940 | 33.01 | 0.918 | 33.00 | 0.925 | 170.73/88.67 | ||
KiT [34] | 22’CVPR | 30.26 | 0.904 | 30.47 | 0.896 | 36.65 | 0.970 | 33.85 | 0.941 | 32.81 | 0.918 | 32.81 | 0.926 | 43.08/20.60 | ||
MAXIM-2S [32] | 22’CVPR | 31.17 | 0.922 | 30.81 | 0.903 | 38.06 | 0.976 | 33.80 | 0.943 | 32.37 | 0.921 | 33.24 | 0.933 | 216.00/14.10 | ||
Restormer [10] | 22’CVPR | 32.00 | 0.923 | 31.46 | 0.904 | 38.99 | 0.978 | 34.18 | 0.944 | 33.19 | 0.926 | 33.96 | 0.935 | 140.99/26.13 | ||
SFNet [78] | 23’ICLR | 31.47 | 0.919 | 31.90 | 0.903 | 38.21 | 0.974 | 33.69 | 0.937 | 32.55 | 0.911 | 33.56 | 0.929 | 125.43/13.27 | ||
MambaIR [79] | 24’ECCV | 32.11 | 0.924 | 32.03 | 0.905 | 38.46 | 0.976 | 34.15 | 0.945 | 33.53 | 0.930 | 34.06 | 0.936 | 229.80/26.78 | ||
@@
Method | Venue | Raindrop-A[57] | Raindrop-B[57] |
PSNR | SSIM | PSNR | SSIM | |||||
pix2pix [80] | 17’CVPR | 28.02 | 0.855 | 23.50 | 0.715 | 18.16/54.40 | ||
AttentiveGAN [57] | 18’CVPR | 31.55 | 0.902 | 24.92 | 0.809 | 89.67/6.24 | ||
DuRN [4] | 19’CVPR | 31.24 | 0.926 | 25.32 | 0.817 | 89.39/12.86 | ||
IDT[81] | 22’TPAMI | 31.87 | 0.931 | - | - | 61.90/16.00 | ||
MAXIM-2S[32] | 22’CVPR | 31.87 | 0.935 | 25.74 | 0.827 | 216.00/14.10 | ||
UDR-S2Former[82] | 23’CVPR | 32.64 | 0.943 | - | - | 21.58/8.53 | ||
@@ IV-B Image Deraining Results
Method | Venue | GoPro[58] | HIDE[59] | RealBlur-R[60] | RealBlur-J[60] |
PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |||||
Nah et al. [58] | 17’CVPR | 29.08 | 0.914 | 25.73 | 0.874 | 32.51 | 0.841 | 27.87 | 0.827 | 336.00/11.70 | ||
DeblurGAN [26] | 18’CVPR | 28.70 | 0.858 | 24.51 | 0.871 | 33.79 | 0.903 | 27.97 | 0.834 | 16.96/14.00 | ||
DeblurGAN-v2 [83] | 19’ICCV | 29.55 | 0.934 | 26.61 | 0.875 | 35.26 | 0.944 | 28.70 | 0.866 | 42.00/60.93 | ||
DMPHN [84] | 19’CVPR | 31.20 | 0.940 | 29.09 | 0.924 | 35.70 | 0.948 | 28.42 | 0.860 | 235.00/21.70 | ||
DBGAN [85] | 20’CVPR | 31.10 | 0.942 | 28.94 | 0.915 | 33.78 | 0.909 | 24.93 | 0.745 | 759.85/11.59 | ||
MTRNN [86] | 20’ECCV | 31.15 | 0.945 | 29.15 | 0.918 | 35.79 | 0.951 | 28.44 | 0.862 | 13.72/2.64 | ||
MIMO-UNet+ [27] | 21’ICCV | 32.45 | 0.957 | 29.99 | 0.930 | 35.54 | 0.947 | 27.63 | 0.837 | 154.41/16.11 | ||
Stripformer [87] | 22’ECCV | 33.08 | 0.962 | 31.03 | 0.940 | - | - | - | - | 177.43/19.71 | ||
DDANet [88] | 23’AAAI | 33.07 | 0.962 | 30.64 | 0.937 | 35.81 | 0.951 | - | - | 153.51/16.18 | ||
MPRNet [30] | 21’CVPR | 32.66 | 0.959 | 30.96 | 0.939 | 35.99 | 0.952 | 28.70 | 0.873 | 777.01/20.13 | ||
NAFNet [52] | 22’ECCV | 33.08 | 0.963 | 31.22 | 0.943 | 36.14 | 0.955 | 28.43 | 0.860 | 63.33/67.89 | ||
Restormer [10] | 22’CVPR | 32.92 | 0.961 | 31.22 | 0.942 | 36.19 | 0.957 | 28.96 | 0.879 | 140.99/26.11 | ||
MambaIR [79] | 24’ECCV | 33.27 | 0.963 | 31.11 | 0.941 | 36.15 | 0.955 | 28.99 | 0.878 | 229.80/26.78 | ||
@@ IV-C Image Deblurring Results
Dataset | Evaluation | DnCNN | RIDNet | ADFNet | MIRNet | MPRNet | Uformer-B | MAXIM-3S | Restormer | NAFNet | MambaIR | MDDA-former | ||||||||||||||||||||||||||
[2] | [90] | [91] | [92] | [30] | [11] | [32] | [10] | [52] | [79] | |||||||||||||||||||||||||||||
Venue | 17’TIP | 19’ICCV | 23’AAAI | 20’ECCV | 21’CVPR | 22’CVPR | 22’CVPR | 22’CVPR | 22’ECCV | 24’ECCV | - | |||||||||||||||||||||||||||
@@ IV-D Image Denoising Results
Method | Venue | SOTS-Indoor[63] | SOTS-Outdoor[63] | SOTS-Mix[64] | Haze4K[65] |
PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |||||
DCP [93] | 09’CVPR | 16.62 | 0.818 | 19.14 | 0.861 | 17.88 | 0.816 | 14.01 | 0.760 | - | ||
DehazeNet [20] | 16’TIP | 19.82 | 0.821 | 24.75 | 0.927 | 21.02 | 0.870 | 19.12 | 0.840 | 0.58/0.009 | ||
MSCNN [94] | 16’ECCV | 19.84 | 0.833 | 22.06 | 0.908 | 20.31 | 0.863 | 14.01 | 0.510 | 0.53/0.008 | ||
AOD-Net [95] | 17’ICCV | 20.51 | 0.816 | 24.14 | 0.920 | 20.27 | 0.855 | 17.15 | 0.830 | 0.12/0.002 | ||
GridDehazeNet [96] | 19’ICCV | 32.16 | 0.984 | 30.86 | 0.982 | 25.86 | 0.944 | 23.29 | 0.930 | 21.49/0.96 | ||
MSBDN [97] | 20’CVPR | 33.67 | 0.985 | 33.48 | 0.982 | 28.56 | 0.966 | 22.99 | 0.850 | 41.54/31.35 | ||
FFA-Net [6] | 20’AAAI | 36.39 | 0.989 | 33.57 | 0.984 | 29.96 | 0.973 | 26.96 | 0.950 | 287.80/4.46 | ||
AECR-Net [21] | 21’CVPR | 37.17 | 0.990 | - | - | 28.52 | 0.964 | - | - | 52.20/2.61 | ||
DeHamer [98] | 22’CVPR | 36.63 | 0.988 | 35.18 | 0.986 | - | - | - | - | 48.93/132.5 | ||
DehazeFormer-M [99] | 23’TIP | 38.46 | 0.994 | 34.29 | 0.983 | 30.89 | 0.977 | - | - | 48.64/4.63 | ||
DehazeFormer-L [99] | 23’TIP | 40.05 | 0.996 | - | - | - | - | - | - | 279.70/25.44 | ||
MITNet [100] | 23’ACM MM | 40.23 | 0.992 | 35.18 | 0.988 | - | - | - | - | 16.42/2.73 | ||
MAXIM-2S [32] | 22’CVPR | 38.11 | 0.991 | 34.19 | 0.985 | - | - | - | - | 216.00/14.10 | ||
MambaIR [79] | 24’ECCV | 40.58 | 0.995 | 35.41 | 0.986 | 31.24 | 0.979 | 33.17 | 0.987 | 229.80/26.78 | ||
@@
Method | Venue | LOL-v1 [66] | LOL-v2-real [67] |
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||||
Retinex-Net [66] | 18’BMVC | 16.77 | 0.560 | 0.381 | 18.37 | 0.723 | 0.425 | 587.47/0.84 | ||
KinD [101] | 19’ACM MM | 20.86 | 0.790 | 0.170 | 19.74 | 0.761 | 0.362 | 34.99/8.16 | ||
Zero-DCE [102] | 20’CVPR | 14.83 | 0.531 | 0.335 | 14.32 | 0.511 | 0.483 | 5.17/0.08 | ||
KinD++ [103] | 21’IJCV | 21.30 | 0.821 | 0.160 | - | - | - | 743.95/8.28 | ||
EnlightenGAN [104] | 21’TIP | 17.48 | 0.650 | 0.322 | 18.23 | 0.617 | 0.421 | 61.01/8.64 | ||
Sparse [67] | 21’TIP | 17.20 | 0.640 | - | 20.06 | 0.816 | - | 53.26/2.33 | ||
DRBN [105] | 21’TIP | 19.55 | 0.746 | 0.316 | 20.13 | 0.820 | 0.304 | 48.61/5.27 | ||
URetinex-Net [106] | 22’CVPR | 21.33 | 0.835 | 0.122 | - | - | - | 56.93/0.40 | ||
UHDFour [107] | 23’ICLR | 23.09 | 0.870 | - | 21.78 | 0.870 | - | -/17.54 | ||
MIRNet [92] | 20’ECCV | 24.14 | 0.830 | - | 20.02 | 0.820 | - | 787.39/31.79 | ||
Restormer [10] | 22’CVPR | 22.43 | 0.823 | 0.141 | 19.94 | 0.827 | 0.292 | 140.99/26.13 | ||
MAXIM-2S [32] | 22’CVPR | 23.43 | 0.863 | - | - | - | - | 216.00/14.10 | ||
MambaIR [79] | 24’ECCV | 24.05 | 0.857 | 0.113 | 22.11 | 0.834 | 0.262 | 229.80/26.78 | ||